OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is -11, the last term is -171,875, and the common ratio is -5? -143,231 144,177 716,144 -36,047

OpenStudy (anonymous):

@ParthKohli would you mind helping me through this?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I don't know where to start? Which formula?

OpenStudy (anonymous):

\[Sn=a \frac{ r^n-1 }{ r-1 }\] n=7 a=-11 r=-7 calculate S7

OpenStudy (anonymous):

S7=-11 - 823544/ -8, and after calculating I get 102932 but this isn't an option :(

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

Am I calculating wrong?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

sorry r=-5

OpenStudy (anonymous):

I got 28645.8333. this isn't right. Can you please walk me through it?

OpenStudy (anonymous):

then r is not =-5 let me calculate

OpenStudy (helder_edwin):

following the formula given to u: a=-11 and r=-5 so \[\large S_7=(-11)\frac{(-5)^7-1}{(-5)-1}=(-11)\frac{-78126}{-6}=-143231 \]

OpenStudy (anonymous):

ahh, thank you <33333

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

so your calculation was wrong.